A 4 X 4 X 4 cubical box contains 64 identical small cubes that exactly fill the box. how many of the small cubes touch a side or the bottom of the box ?
a) 48 b) 52 c) 60 d) 64 e) 80

Answer by checkley77(12844)

The 2 ends=16each or 2*16=32
The 2 sides=8 each or 2*8=16
The bottom=4
32=16+4=52 answer b).